OpenStudy (anonymous):

compute the derivative of 3^x+1/x+2

OpenStudy (anonymous):

So we want to find \(p\) given \(\large e^p=3^x\).

OpenStudy (anonymous):

Does that much make sense?

OpenStudy (anonymous):

We want \(p\) in terms of \(x\) since we don't know \(x\). @abbyhernandez

OpenStudy (anonymous):

How would you solve for \(p\)?

OpenStudy (anonymous):

ok question changed...

OpenStudy (anonymous):

i think it would be a good idea to rewrite it in terms of e. but I'm a bit confused as how to go about that. i have e^ln3(x+1/x+2). not sure how to proceede from here

OpenStudy (anonymous):

is the function \[\LARGE 3^{\frac{x+1}{x+2}} \]

OpenStudy (anonymous):

????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know the chain rule?

OpenStudy (anonymous):

Yes I know the chain rule : ) I am at work and have to log out for a while. i apprecite if you can help. i will log in as soon as i get home.

OpenStudy (dumbcow):

derivative of form a^f(x) \[\frac{d}{dx} a^{f(x)} = \ln(a)*a^{f(x)}*f'(x)\]